OpenStudy (geometry_hater):

The diagonal of a square is 8 square root of 2 units. Which is the perimeter of the square?

OpenStudy (geometry_hater):

8 16 24 32

OpenStudy (slaaibak):

 x^2 + x^2 = (8sqrt2)^2 2x^2=128 x^2=64 x=8 perimeter = 4 * 8 = 32